# Non-toric brane webs, Calabi–Yau 3-folds, and 5d SCFTs

### v1

@@ 4.1. Consistent asymptotic webs of $5$-branes with $7$-branes and 5d SCFTs
In this section, we first define a geometric notion of supersymmetric asymptotic webs, that are expected to define 5d SCFTs. Then, we introduce the more restrictive notion of consistent asymptotic webs, and we argue that any 5d SCFT defined by a supersymmetric web can be obtained from a 5d SCFT defined by a consistent web by adding free hypermultiplets.
@@
As described in §3.3, the log Calabi–Yau surface $(Y,D)$ is obtained from a toric surface $(\overline{Y},\overline{D})$ by interior blow-ups of points $y_{e,i}$ on $\overline{D}$, with exceptional curves $E_{e,i}$. Moreover, by Equation (3.3), we have $L=p^{\star}\overline{L}\otimes\mathcal{O}(-\sum_{e,i}a_{e,i}E_{e,i})$, where $a_{e,i}\in\mathbb{Z}_{\geq 1}$ is the number of 5-branes ending on the 7-brane $x_{e,i}$, and $p:\overline{Y}\rightarrow Y$ is the blow-up map. In particular, if $C$ is a smooth curve in $|L|$ that does not contain any of the exceptional curves $E_{e,i}$, we have $C\cdot E_{e,i}=a_{e,i}$, and the image $\overline{C}:=p(C)$ of $C$ in $\overline{Y}$ is a curve in the linear system $|\overline{L}|$ passing through the point $y_{e,i}$ with multiplicity $a_{e,i}$. In particular, $y_{e,i}$ is a singular point of $\overline{C}$ if $a_{e,i}>1$, and $p|_{C}:C\rightarrow\overline{C}$ is a resolution of the singularities of $\overline{C}$ that separate the $a_{e,i}$ branches of $\overline{C}$ passing through $y_{e,i}$ – see Figure 4.1.
@@
In the physics literature, the proposed Seiberg–Witten curves for the compactification on $S^{1}$ of a 5d SCFT defined by a web of 5-branes with 7-branes are the curves $\overline{C}^{\circ}=\overline{C}\cap(\mathbb{C}^{\star})^{2}$. For instance, the condition that $y_{e,i}$ is a multiple point of the curve $\overline{C}$ with multiplicity $a_{e,i}$ is equivalent to the divisibility conditions on Laurent polynomials proposed in [67, (2.15)] to characterize Seiberg–Witten curves. Instead, we propose to view the curve $C^{\circ}=C\cap U$ in the open Calabi–Yau surface $U$ as the proper description of the Seiberg–Witten curve.
[…]
Figure 4.1. A curve $C\in|L|$ in the log Calabi–Yau surface $(Y,D)$, and the corresponding curve $\overline{C}\in|\overline{L}|$ in the toric surface $(\overline{Y},\overline{D})$.
In what follows, we compare the notion of supersymmetric asymptotic web given by Definition 4.1 with supersymmetric conditions usually formulated in the physics literature in terms of Hanany–Witten moves on webs of 5-branes with 7-branes. Recall from §3.1 that Hanany–Witten moves on configurations of $7$-branes are defined by pushing a focus-focus singularity across the origin, and rotating the adjacent cut clockwise, while changing the positions of all other $7$-branes by a shear as in Equation (3.1). Hanany–Witten moves on webs of $5$-branes with $7$-branes change in addition the configuration of $5$-branes as follows. Consider the Hanany–Witten move obtained by moving the $(p_{e},q_{e})$ $7$-brane $x_{e,i}$ across the origin, for a leg $e\in L(W^{\mathrm{asym}})$ and some $1\leq i\leq n_{e}$. The 7-brane $x_{e,i}$ becomes a $(-p_{e},-q_{e})$ $7$-brane $x^{\prime}_{e,i}$, and the other 7-branes are changed according to the shear transformation defined by Equation (3.1). We apply the same shear transformation to all the $5$-branes that do not end at $x_{e,i}$. Moreover, the $5$-branes that ended in $x_{e,i}$ disappear, and the number $a_{e,i}^{\prime}$ of new 5-branes passing through the origin and ending on $x_{e,i}^{\prime}$ is determined by imposing the balancing condition of the 5-brane charge at the origin. The balancing condition uniquely determines $a_{e,i}^{\prime}$ as an integer, but this integer is not necessarily positive. Denoting by $w_{-e}\in\mathbb{Z}_{\geq 0}$ the weight of the leg of direction $(-p_{e},-q_{e})$ in $\overline{W}^{\mathrm{asym}}$, so that the weight of the edge connecting the origin to $x_{e,i}^{\prime}$ is $w_{-e}+a_{e,i}^{\prime}$, see Figure 4.2, we are in one of the three following situations:
@@
    $-w_{-e}\leq a_{e,i}^{\prime}\leq 0\,.$ (4.2)
    In this situation, the 7-brane $x_{e,i}^{\prime}$ can be viewed as the endpoint of $-a_{e,i}^{\prime}$ 5-branes which come from infinity, or ending on another 7-brane, see Figure 4.3. These 5-branes are not connected to the part of the web of 5-branes containing the origin, and so can be pushed to infinity together with the 7-brane $x_{e,i}^{\prime}$.
@@
[…]
Figure 4.2. Hanany–Witten move for a web of 5-branes with 7-branes.
In the physics literature, Case I is interpreted as a sign that the web is not supersymmetric since an edge with a negative multiplicity can be viewed as a supersymmetry-breaking anti-5-brane. We show in Lemma 4.3 below that this interpretation is compatible with our geometric Definition 4.1 of a supersymmetric web. To do that, we first explain that the above combinatorial description of Hanany–Witten moves at the level of webs of 5-branes with 7-branes has a natural geometric interpretation in terms of elementary cluster transformations of log Calabi–Yau surfaces, as discussed in §3.1.3 in the case of configurations of 7-branes.
Recall from §3.3 that the log Calabi–Yau surface $(Y,D)$ is a non-toric blow-up of a toric surface $(\overline{Y},\overline{D})$ with fan given by the support of the web of 5-branes $\overline{W}^{\mathrm{asym}}$. Up to replacing $(\overline{Y},\overline{D})$ by a toric blow-up, one can assume that the fan of $(\overline{Y},\overline{D})$ contains both rays $\mathbb{R}_{\geq 0}(p_{e},q_{e})$, and $\mathbb{R}_{\geq 0}(-p_{e},-q_{e})$. Then, as in §3.1.3, the projection $\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}/\mathbb{R}(p_{e},q_{e})=\mathbb{R}$ induces a toric morphism $\nu:\overline{Y}\rightarrow\mathbb{P}^{1}$, which is generically a $\mathbb{P}^{1}$-fibration. Moreover, the irreducible components $D_{e}$ and $D_{e}^{\prime}$ of $\overline{D}$ corresponding respectively to the rays $\mathbb{R}_{\geq 0}(p_{e},q_{e})$ and $\mathbb{R}_{\geq 0}(-p_{i},-q_{i})$ of the fan, are sections of $\nu$. The construction of $Y$ from $\overline{Y}$ involves blowing up the point $y_{e,i}$ on $D_{e}$, which creates an exceptional $(-1)$-curve $E_{e,i}$ in $Y$. The strict transform of the $\mathbb{P}^{1}$-fiber of $\nu$ passing through $x_{e,i}$ is also a $(-1)$-curve $E_{e,i}^{\prime}$. Hence, one can construct a new toric surface $\overline{Y}^{\prime}$ from $Y$ by contracting all the exceptional curves coming from blowing up the points $x_{f,j}$ with $(f,j)\neq(e,i)$, and the $(-1)$-curve $E_{e,i}^{\prime}$. Geometrically, the Hanany–Witten move from $x_{e,i}$ to $x_{e,i}^{\prime}$ encodes the passage from describing $Y$ as a blow-up of the toric surface $\overline{Y}$ to describing $Y$ as a blow-up of the toric surface $\overline{Y}^{\prime}$. In particular, the weight $w_{-e}$ in $W^{\mathrm{asym}}$ is given by the intersection number
$w_{-e}=L\cdot D_{e}^{\prime}\,,$ (4.4)
@@
Since $W^{\mathrm{asym}}$ is supersymmetric, it follows from Definition 4.1 that there exists a curve $C$ in $|L|$ not containing any 0-dimensional stratum of $D$. Keeping the notation used in the discussion of the geometric interpretation of Hanany–Witten moves, let $E_{e,i}^{\prime}$ be the $(-1)$-curve corresponding to the 7-brane $x_{e,i}^{\prime}$. Denote by $C^{\prime}$ the union of irreducible components of $C$ not containing $E_{e,i}^{\prime}$, so that $C=C^{\prime}+wE_{e,i}^{\prime}$ for some $w\in\mathbb{Z}_{\geq 0}$. By Equations (4.5) and (4.4), we have
@@
    For every curve $E$ in $Y$ whose strict transform in a corner blow-up of $Y$ is either an interior $(-1)$-curve or an internal $(-2)$-curve, we have $L\cdot E>0$.
@@
The fact that ii) implies i) follows from the above geometric interpretation of Cases I-III in (4.1)–(4.3) for Hanany–Witten moves. Hence, it remains to show that i) implies ii). Let $E$ be a curve in $Y$ whose strict transform in a corner blow-up of $Y$ is an irreducible $(-1)$-curve. By [55, Proposition 3.27] (see also [21, Theorem 1]), any two toric models of a log Calabi–Yau surface are connected by a sequence of corner blow-ups/blow-downs and elementary cluster transformations. Hence, there exists a sequence of Hanany–Witten moves on $W^{\mathrm{asym}}$ such that $E$ is the $(-1)$-curve $E_{e,i}^{\prime}$ corresponding to the pushed 7-brane $x_{e,i}^{\prime}$. The inequality (4.3) for this sequence of Hanany–Witten moves implies that $L\cdot E>0$. One can argue similarly for an internal $(-2)$-curve and this concludes the proof. ∎
@@
    For every curve $E$ in $Y$ whose strict transform in a corner blow-up of $Y$ is an interior $(-1)$-curve or an internal $(-2)$-curve, we have $L\cdot E>0$.
@@
    A 7-brane is sent to infinity as on the left of Figure 4.3. In this case, there exists an interior $(-1)$-curve $E$ such that $L\cdot E=0$ and that can be contracted to a point.
@@
    A 5-brane attached to a 7-brane is sent to infinity as in the middle of Figure 4.3. In this case, there is an interior $(-1)$-curve $E$ such as a multiple $kE$ is a connected component of $C$, and so $C$ can be replaced by $C\setminus kE$.
@@
    A 5-brane connected to two 7-branes is sent to infinity as on the right of Figure 4.3. In this case, there is an internal $(-2)$-curve $E$ such as a multiple $kE$ is a connected component of $C$, and so $C$ can be replaced by $C\setminus kE$.
Iterating these operations, we finally obtain a log Calabi–Yau surface with line bundle $(Y^{\prime},D^{\prime},L^{\prime})$ corresponding to a consistent asymptotic web. In particular, the curve $C\in|L|$ is isomorphic to the disjoint union of a curve isomorphic to a curve $C^{\prime}\in|L^{\prime}|$ and of finitely many disjoint multiple interior $(-1)$-curves and internal $(-2)$-curves in $Y$. Physically, every 7-brane or 5-brane sent to infinity corresponds to a free hypermultiplet, as described in [19, Appendix A.3]. Hence, the 5d SCFT defined by a supersymmetric asymptotic web of 5-branes with 7-branes can always be obtained from the 5d SCFT defined by a consistent web by adding free hypermultiplets. Therefore, in what follows, we focus our study on consistent webs.
[…]
Figure 4.3. Possible results of a Hanany–Witten move in Case II as in Equation (4.2).

### v2

@@ 4.1. Consistent asymptotic webs of $5$-branes with $7$-branes and 5d SCFTs
In this section, we first define a geometric notion of supersymmetric asymptotic webs, that are expected to define 5d SCFTs. Then, we introduce the more restrictive notion of consistent asymptotic webs, and we argue that any 5d SCFT defined by a supersymmetric web can be obtained from a 5d SCFT defined by a consistent web by adding free hypermultiplets. In particular, restricting attention to consistent webs entails no loss of generality when constructing non-trivial 5d SCFTs.
@@
As described in §3.3, the log Calabi–Yau surface $(Y,D)$ is obtained from a toric surface $(\overline{Y},\overline{D})$ by interior blow-ups of points $y_{e,i}$ on $\overline{D}$, with exceptional curves $E_{e,i}$. Moreover, by Equation (3.3), we have $L=p^{\star}\overline{L}\otimes\mathcal{O}(-\sum_{e,i}a_{e,i}E_{e,i})$, where $a_{e,i}\in\mathbb{Z}_{\geq 1}$ is the number of 5-branes ending on the 7-brane $x_{e,i}$, and $p:\overline{Y}\rightarrow Y$ is the blow-up map. In particular, if $C$ is a smooth curve in $|L|$ that does not contain any of the exceptional curves $E_{e,i}$, we have $C\cdot E_{e,i}=a_{e,i}$, and the image $\overline{C}:=p(C)$ of $C$ in $\overline{Y}$ is a curve in the linear system $|\overline{L}|$ passing through the point $y_{e,i}$ with multiplicity $a_{e,i}$. In particular, $y_{e,i}$ is a singular point of $\overline{C}$ if $a_{e,i}>1$, and $p|_{C}:C\rightarrow\overline{C}$ is a resolution of the singularities of $\overline{C}$ that separate the $a_{e,i}$ branches of $\overline{C}$ passing through $y_{e,i}$ – see Figure 8.
@@
In the physics literature, the proposed Seiberg–Witten curves for the compactification on $S^{1}$ of a 5d SCFT defined by a web of 5-branes with 7-branes are the curves $\overline{C}^{\circ}=\overline{C}\cap(\mathbb{C}^{\star})^{2}$. For instance, the condition that $y_{e,i}$ is a multiple point of the curve $\overline{C}$ with multiplicity $a_{e,i}$ is equivalent to the divisibility conditions on Laurent polynomials proposed in [70, (2.15)] to characterize Seiberg–Witten curves. Instead, we propose to view the curve $C^{\circ}=C\cap U$ in the open Calabi–Yau surface $U$ as the proper description of the Seiberg–Witten curve.
Figure 8. A curve $C\in|L|$ in the log Calabi–Yau surface $(Y,D)$, and the corresponding curve $\overline{C}\in|\overline{L}|$ in the toric surface $(\overline{Y},\overline{D})$.
In what follows, we compare the notion of supersymmetric asymptotic web given by Definition 4.1 with supersymmetric conditions usually formulated in the physics literature in terms of Hanany–Witten moves on webs of 5-branes with 7-branes. Recall from §3.1 that Hanany–Witten moves on configurations of $7$-branes are defined by pushing a focus-focus singularity across the origin, and rotating the adjacent cut clockwise, while changing the positions of all other $7$-branes by a shear as in Equation (3.1). Hanany–Witten moves on webs of $5$-branes with $7$-branes change in addition the configuration of $5$-branes as follows. Consider the Hanany–Witten move obtained by moving the $(p_{e},q_{e})$ 77-brane $x_{e,i}$ across the origin, for a leg $e\in L(W^{\mathrm{asym}})$ and some $1\leq i\leq n_{e}$. The 7-brane $x_{e,i}$ becomes a $(-p_{e},-q_{e})$ 77-brane $x^{\prime}_{e,i}$, and the other 7-branes are changed according to the shear transformation defined by Equation (3.1). We apply the same shear transformation to all the $5$-branes that do not end at $x_{e,i}$. Moreover, the $5$-branes that ended in $x_{e,i}$ disappear, and the number $a_{e,i}^{\prime}$ of new 5-branes passing through the origin and ending on $x_{e,i}^{\prime}$ is determined by imposing the balancing condition of the 5-brane charge at the origin. The balancing condition uniquely determines $a_{e,i}^{\prime}$ as an integer, but this integer is not necessarily positive. Denoting by $w_{-e}\in\mathbb{Z}_{\geq 0}$ the weight of the leg of direction $(-p_{e},-q_{e})$ in $\overline{W}^{\mathrm{asym}}$, so that the weight of the edge connecting the origin to $x_{e,i}^{\prime}$ is $w_{-e}+a_{e,i}^{\prime}$, see Figure 9, we are in one of the three following situations:
@@
    $w_{-e}\leq a_{e,i}^{\prime}\leq 0\,.$ (4.2)
    In this situation, the 7-brane $x_{e,i}^{\prime}$ can be viewed as the endpoint of $a_{e,i}^{\prime}$ 5-branes which come from infinity, or ending on another 7-brane, see Figure 10. These 5-branes are not connected to the part of the web of 5-branes containing the origin, and so can be pushed to infinity together with the 7-brane $x_{e,i}^{\prime}$.
@@
Figure 9. Hanany–Witten move for a web of 5-branes with 7-branes.
In the physics literature, Case I is interpreted as evidence that the web is not supersymmetric since an edge with a negative multiplicity can be viewed as a supersymmetry-breaking anti-5-brane. We show in Lemma 4.3 below that this interpretation is compatible with our geometric Definition 4.1 of a supersymmetric web. To do that, we first explain that the above combinatorial description of Hanany–Witten moves at the level of webs of 5-branes with 7-branes has a natural geometric interpretation in terms of elementary cluster transformations of log Calabi–Yau surfaces, as discussed in §3.1.3 in the case of configurations of 7-branes.
Recall from §3.3 that the log Calabi–Yau surface $(Y,D)$ is a non-toric blow-up of a toric surface $(\overline{Y},\overline{D})$ with fan given by the support of the web of 5-branes $\overline{W}^{\mathrm{asym}}$. Up to replacing $(\overline{Y},\overline{D})$ by a toric blow-up, one can assume that the fan of $(\overline{Y},\overline{D})$ contains both rays $\mathbb{R}_{\geq 0}(p_{e},q_{e})$, and $\mathbb{R}_{\geq 0}(-p_{e},-q_{e})$. Then, as in §3.1.3, the projection $\mathbb{R}^{2}\rightarrow\mathbb{R}^{2}/\mathbb{R}(p_{e},q_{e})=\mathbb{R}$ induces a toric morphism $\nu:\overline{Y}\rightarrow\mathbb{P}^{1}$, which is generically a $\mathbb{P}^{1}$-fibration. Moreover, the irreducible components $\overline{D}_{e}$ and $\overline{D}_{e}^{\prime}$ of $\overline{D}$ corresponding respectively to the rays $\mathbb{R}_{\geq 0}(p_{e},q_{e})$ and $\mathbb{R}_{\geq 0}(-p_{i},-q_{i})$ of the fan, are sections of $\nu$. The construction of $Y$ from $\overline{Y}$ involves blowing up the point $y_{e,i}$ on $\overline{D}_{e}$, which creates an exceptional (1)(-1)-curve $E_{e,i}$ in $Y$. The strict transform of the $\mathbb{P}^{1}$-fiber of $\nu$ passing through $y_{e,i}$ is also a (1)(-1)-curve $E_{e,i}^{\prime}$, which intersects the strict transform $D_{e}^{\prime}$ of $\overline{D}_{e}^{\prime}$. Hence, one can construct a new toric surface $\overline{Y}^{\prime}$ from $Y$ by contracting all the exceptional curves coming from blowing up the points $y_{f,j}$ with $(f,j)\neq(e,i)$, and the (1)(-1)-curve $E_{e,i}^{\prime}$. Geometrically, the Hanany–Witten move from $x_{e,i}$ to $x_{e,i}^{\prime}$ encodes the passage from describing $Y$ as a blow-up of the toric surface $\overline{Y}$ to describing $Y$ as a blow-up of the toric surface $\overline{Y}^{\prime}$. In particular, the weight $w_{-e}$ in $W^{\mathrm{asym}}$ is given by the intersection number
$w_{-e}=L\cdot D_{e}^{\prime}\,,\,$, (4.4)
@@
Since $W^{\mathrm{asym}}$ is supersymmetric, it follows from Definition 4.1 that there exists a curve $C$ in $|L|$ not containing any 0-dimensional stratum of $D$. Keeping the notation used in the discussion of the geometric interpretation of Hanany–Witten moves, let $E_{e,i}^{\prime}$ be the (1)(-1)-curve corresponding to the 7-brane $x_{e,i}^{\prime}$. Denote by $C^{\prime}$ the union of irreducible components of $C$ not containing $E_{e,i}^{\prime}$, so that $C=C^{\prime}+wE_{e,i}^{\prime}$ for some $w\in\mathbb{Z}_{\geq 0}$. By Equations (4.5) and (4.4), we have
@@
    For every curve $E$ in $Y$ whose strict transform in a corner blow-up of $Y$ is either an interior (1)(-1)-curve or an internal (2)(-2)-curve, we have $L\cdot E>0$.
@@
The fact that ii) implies i) follows from the above geometric interpretation of Cases I-III in (4.1)–(4.3) for Hanany–Witten moves. Hence, it remains to show that i) implies ii). Let $E$ be a curve in $Y$ whose strict transform in a corner blow-up of $Y$ is an irreducible (1)(-1)-curve. By [58, Proposition 3.27] (see also [24, Theorem 1]), any two toric models of a log Calabi–Yau surface are connected by a sequence of corner blow-ups/blow-downs and elementary cluster transformations. Hence, there exists a sequence of Hanany–Witten moves on $W^{\mathrm{asym}}$ such that $E$ is the (1)(-1)-curve $E_{e,i}^{\prime}$ corresponding to the pushed 7-brane $x_{e,i}^{\prime}$. The inequality (4.3) for this sequence of Hanany–Witten moves implies that $L\cdot E>0$. One can argue similarly for an internal (2)(-2)-curve and this concludes the proof. ∎
@@
    For every curve $E$ in $Y$ whose strict transform in a corner blow-up of $Y$ is an interior (1)(-1)-curve or an internal (2)(-2)-curve, we have $L\cdot E>0$.
@@
    A 7-brane is sent to infinity as on the left of Figure 10. In this case, there exists an interior (1)(-1)-curve $E$ such that $L\cdot E=0$ and that can be contracted to a point.
@@
    A 5-brane attached to a 7-brane is sent to infinity as in the middle of Figure 10. In this case, there is an interior (1)(-1)-curve $E$ such as a multiple $kE$ is a connected component of $C$, and so $C$ can be replaced by $C\setminus kE$.
@@
    A 5-brane connected to two 7-branes is sent to infinity as on the right of Figure 10. In this case, there is an internal (2)(-2)-curve $E$ such as a multiple $kE$ is a connected component of $C$, and so $C$ can be replaced by $C\setminus kE$.
Each instance of case i) strictly decreases the Picard rank of $Y$, and thus case i) can occur only finitely many times. Cases (ii) and (iii) are also limited in number, since the number of connected components of a curve in the linear system $|L|$ is bounded. Therefore, iterating these operations finitely many times, we finally obtain a log Calabi–Yau surface with line bundle $(Y^{\prime},D^{\prime},L^{\prime})$ corresponding to a consistent asymptotic web. The curve $C\in|L|$ is isomorphic to the disjoint union of a curve isomorphic to a curve $C^{\prime}\in|L^{\prime}|$ and of finitely many disjoint multiple interior (1)(-1)-curves and internal (2)(-2)-curves in $Y$. Physically, every 7-brane or 5-brane sent to infinity corresponds to a free hypermultiplet, as described in [22, Appendix A.3]. Hence, the 5d SCFT defined by a supersymmetric asymptotic web of 5-branes with 7-branes can always be obtained from the 5d SCFT defined by a consistent web by adding free hypermultiplets. Therefore, in what follows, we focus our study on consistent webs.
Figure 10. Possible results of a Hanany–Witten move in Case II as in Equation (4.2).